OpenStudy (anonymous):

Mr whippy decides that he wants to use paper containers for his customers to hold their individual icecream cones. How much paper will he need for one cone if his ice cream cone has a diameter of 6cm, and the slant edge is 15cm long?(the cone does not need a top) Give your answer correct to the nearest cm^2.

OpenStudy (anonymous):

The lateral surface area of a cone is L.A. = pi*r*l where pi = 3.14, r = radius of the cone, and l = slant height of the cone. L.A. = pi*3*15 = 45*pi = 141.37 = 142cm^2